# Controlled Erasure as a Building Block for Universal Thermodynamically-Robust Superconducting Computing

### v1

@@ V.1 Effective CE Protocol
Fig. 5 highlights the important fixed points on the potential immediately before and after the stable (red) and unstable (green) fixed points annihilate in the negative-$x$ half plane. After the annihilation, a CE is possible since the orange, blue, and purple fixed points on the positive-$x$ half plane are preserved. From the perspective detailed in Sec. III, the information stored in the $\mathtt{Y}$ state on the negative-$x$ half plane is erased, while all other information in the potential’s memory states is preserved. In this example, $\Delta U_{00}^{\prime}$ denotes the potential barrier required to vanish for the CE to be carried out. Once this barrier vanishes, $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$ and $\Delta U_{11}^{\prime}$ indicate the barriers that need to be maintained. The goal is to find the circuit parameter values which give these barriers sufficiently large heights, as this further contributes to a successful CE.
@@
Fig. 5 illustrates bifurcation diagrams of the $\varphi_{2}$ coordinate over a given circuit parameter range, while all other parameters are held constant. Motivated by Ozfidan et al. [36], we take $\beta=2.3$. Additionally, the coordinate $\varphi_{2}$ is solely illustrated because $\varphi_{1}$ varies negligibly in comparison.
[…]
@@
The following now describes the qualitative consequences of changing each circuit parameter and the resulting potential in Fig. 5. First, as $\mu$ increases, the depth of the minima representing the states $01$ and $10$ increases. Meanwhile, the minima depths in the $11$ state decrease, leading to $\Delta U_{11}^{\prime}$ decreasing. This indicates that $\mu$ should not take on too large of a value in order to avoid fluctuations over the $\Delta U_{11}^{\prime}$ barrier. Next, if the magnitude of $\varphi_{2\mathrm{xdc}}$ increases, then both $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$ decrease. Consequently, its value should be large enough only to ensure that $\Delta U_{00}^{\prime}$ is eliminated while $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$ are maintained. Changing $\varphi_{1\mathrm{x}}$ adjusts the potential’s tilt with respect to the $\varphi_{1}$ axis. If $\varphi_{1\mathrm{x}}$ takes on too large of a value, then the barrier $\Delta U_{01}^{\prime}$ will be undesirably eliminated. However, its value must be large enough to ensure that $\Delta U_{00}^{\prime}$ is eliminated. Similarly, tuning $\varphi_{2\mathrm{x}}$ changes the tilt of the potential with respect to the $\varphi_{2}$ axis. If $\varphi_{2\mathrm{x}}$ is too large, then the blue (purple) stable (unstable) fixed points can annihilate, which causes information to be erased in the region where it should be preserved.
@@
Obtaining effective circuit parameter ranges involves determining two particular values: (i) the value resulting in $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, and $\Delta U_{11}^{\prime}$ having the same height and (ii) the value which results in one barrier falling below the proxy line. The first (second) criterion serves as the lower (upper) ends of the range. Ultimately, there are clear tradeoffs that depend on the value of a given circuit parameter. Finding the set of parameters permitting control over which information is erased and preserved requires a balancing act: A careful process of selecting parameter values that are large enough to erase the desired information, but not so large that the information originally planned to be preserved ends up being erased.
@@
Figure 5: Circuit parameter ranges that eliminate the dimensionless potential barriers $\Delta U_{00}^{\prime}$ and maintain $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$. In the insets, light blue, light green, and light red correspond to the barriers $\Delta U_{11}^{\prime}$, $\Delta U_{10}^{\prime}$ and $\Delta U_{01}^{\prime}$, respectively. The barrier $\Delta U_{01}^{\prime}$ serves as the most conservative one for the $01$ state during the CE. Last, the dotted line corresponds to a thermal energy barrier of $50\;k_{B}T$ at $T=100$ mK. (a) Specific fixed points (colored diamonds) within the potential energy landscape, as well as corresponding energy barriers, are utilized to investigate the CE in Fig. 2. (Left) Prior to the fixed point annihilation in the third quadrant. (Right) After fixed point annihilation that causes $\Delta U_{00}^{\prime}$ to vanish. Once all information stored in the $00$ state is erased into the $01$ state and the potential is subsequently brought back to its original configuration, the CE is completed. (b) Circuit parameter ranges for achieving the CE in Fig. 2, which can be accomplished by the annihilation of the green and red fixed points while maintaining the yellow and blue fixed points. Each circuit parameter range holds all other nonzero circuit parameters at the respective constant values: $\mu=0.06$, $\varphi_{2\mathrm{xdc}}=1.79$, $\varphi_{1\mathrm{x}}=0.59$, and $\varphi_{2\mathrm{x}}=0.10$. These circuit parameters lead to effective CE protocols and thereby robust NAND gates.
@@ V.2 Control Erase Parameter Selection
We can now detail which circuit parameters—including their respective signs and magnitudes—are deployed for any of the eight CE protocols of interest. By leveraging the relationship between the potential’s dynamics and the circuit’s parameters, and the CE notation introduced in Sec. III, all eight CE protocols in Fig. 8 can be implemented in a systematic way. The following explains this approach, summarizing the results in Table 2.
@@
Here, $[P]$ is the Iverson bracket of the statement $P$; $[P]$ evaluates to 1 if $P$ is true, or 00 if $P$ is false. This notation is useful when discussing how the sign of a circuit parameter is determined. Let’s review the four parameters involved in each CE protocol:
@@
For these parameters, $i,j\in\{1,2\}$ with $i\neq j$. To determine which $\varphi_{i\mathrm{xdc}}$ is utilized for a particular CE, recall that we specify a CE with three parameters: $\rho$, $\sigma$, and $A$. If $\rho=\mathtt{X}\;(\mathtt{Y})$, then $i,j=2,1\;(1,2)$. As seen from Eq. (3), $\varphi_{i\mathrm{xdc}}$ appears only in the $\mathrm{cosine}$ function: Since this function is even, the sign of $\varphi_{i\mathrm{xdc}}$ is not relevant. Recall from Secs. III and V.1 that $\varphi_{i\mathrm{xdc}}$ lowers energy barriers on both sides of the $i$th axis, but the CE requires one barrier to be maintained. Thus, we need a control parameter to offset this barrier drop so that it applies to only one half of the $\varphi_{1}$$\varphi_{2}$ plane, while maintaining the other barrier. The parameter deploying this barrier offset is $\varphi_{j\mathrm{x}}$, and it’s sign is given by $\llbracket\sigma\neq\rho\rrbracket$.
Next, $\mu$ biases the potential such that the minima along one of the diagonals of the $\varphi_{1}$$\varphi_{2}$ plane  are energetically favorable compared to those on the other diagonal. Its sign determines which well becomes the target for the erasure, and is given by $\llbracket[\sigma=\rho]\ \mathrm{XOR}\ A\rrbracket$. Recall that $A\in\{0,1\}$ is the state to which the control bit is erased. Since $\mu$ acts along the diagonals of the plane, it has the unwanted effect of making both wells along a given diagonal more energetically favorable. This causes asymmetry between two wells that are supposed to be storing information, resulting unwanted information leakage. Fortunately, the parameter $\varphi_{i\mathrm{x}}$ can be used to compensate this effect by tilting the potential to offset this unintentional information leakage. The sign of this parameter is found through $\llbracket A\rrbracket$.
@@ V.3 CE Robustness
As this information is erased, other information is being stored in the $01$, $10$, and $11$ states due to the barriers $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, and $\Delta U_{11}^{\prime}$ being maintained. However, we see in Fig. 5 that as a circuit parameter increases, at least one barrier’s height decreases. The microstates stored in these decreasing barriers are most susceptible to fluctuating out of their respective memory states. Due to this, the timescale that characterizes how long these states reliably store information is in competition with $t_{\mathrm{CE}}$. This type of timescale is known as a dwell time [16, 18, 37], which is denoted $t_{\mathrm{d}}$. Implicitly from Fig. 5, there are three dwell times that contribute to CE performance—those being the dwell times corresponding to the states $01$, $10$, and $11$. The minimum of these three dwell times ultimately dictates CE failure.
@@
If the length of time that information can be reliably stored is much longer than the timescale of a CE protocol, there is a low probability of error due to an unwanted thermally activated transition. To quantify this, we take the ratio $t_{\mathrm{d}}/t_{\mathrm{CE}}$ as a measure of robustness.
@@
To calculate $t_{\mathrm{CE}}$, we approximate the information in the $00$ state as a single damped particle travelling down a quadratic potential from the point where the fixed points initially annihilate; the derivation is included in App. A. For $t_{\mathrm{d}}$, since there are three dwell times that affect CE performance, we choose the minimum of these dwell times to characterize CE robustness. This provides a more conservative estimate of a particular CE’s performance; refer to App. B for details. From here, we note that the dwell time is exponentially proportional to $L$ and $T$. This has a dominant effect on the robustness ratio $t_{\mathrm{d}}/t_{\mathrm{CE}}$. Therefore, studying the robustness as a function of $L$ and $T$ allows for interpreting the performance of a CE across the MQP and QFP literature.
@@
From Sec. V.1, selecting parameter values at the lower end of the effective parameter range yields a more robust CE: We found that $\mu=0.06$, $\varphi_{i\mathrm{xdc}}=1.79$, $\varphi_{i\mathrm{x}}=0.61$, and $\varphi_{j\mathrm{x}}=0.10$. These will be known as the effective parameter values. With these, Fig. 6 displays the CE robustness for different values of $L$ and $T$. Since the MQP (QFP) literature employ SQUIDs for different purposes—thereby leading to differing ranges of $L$ and $T$—we separate the possible value ranges into the left (right) figures. The relevant MQP literature [38, 17, 16, 18, 32, 19, 33, 27, 35, 29] corresponds to the left figure, which has a general temperature range of $[100,500]\;\text{mK}$, while the inductance values range from $[140,300]\;\text{pH}$. An upper value of $\mathcal{O}(10^{48})$ CEs per device is obtained with the inductance value of $L=140\;\text{pH}$ found from Saira et al. [27] at $T=100\;\text{mK}$.
@@
Now, suppose all circuit parameters were initialized to zero, so that the potential matches Fig. 1. Then, tune all circuit parameters infinitely fast to their respective final values. Due to this, the information stored in the $00$ state is now travelling from the location determined by zero-valued circuit parameters, as opposed to the point of annihilation. Considering this scenario provides a more conservative estimate for the ratio $t_{\mathrm{d}}/t_{\mathrm{CE}}$, as the information from the $00$ state must travel a greater distance to the $01$ state, thereby resulting in a larger $t_{\mathrm{CE}}$. However, this larger $t_{\mathrm{CE}}$ will be no more than twice the value of the most conservative estimate of $t_{\mathrm{CE}}$, leading to a reduction of the robustness of order one compared to that shown in Fig. 6.
@@ V.4 NAND Gate Application
Executing successive CE protocols allows for performing a NAND gate. The CE has two input bits and two output bits as seen from Fig. 3, while the NAND gate has two inputs but only one output bit. When the number of logical output bits is less than the number of embedded output bits in a computational system, there is added flexibility when choosing which bit to read as the output for a computation. For example, we can specify $\mathtt{X}^{\prime}=\text{NAND}[(\mathtt{X},\mathtt{Y})]$ without prescribing the $\mathtt{Y}^{\prime}$ computation. Thus, we define two kinds of NAND gates: a partial NAND, written as $P_{\rho^{\prime}}[(\mathtt{X},\mathtt{Y})]$, which only requires the state $\rho^{\prime}$ to yield the desired output, and a complete NAND, denoted as $C[(\mathtt{X},\mathtt{Y})]$, which requires that $\mathtt{X}^{\prime}=\mathtt{Y}^{\prime}=\text{NAND}[(\mathtt{X},\mathtt{Y})]$.
@@
Fig. 7 illustrates both kinds of computation. Recall from Fig. 1 that the first, second, third, and fourth quadrants correspond to the memory states $11$, $01$, $00$, and $10$, respectively. In this example, carrying out only steps (1)-(3) executes the partial computation $P_{\mathtt{Y}^{\prime}}$. This can seen by tracking the arrows through the included protocol steps: In step (1), the information in the state $10$ is erased into the $00$ state; in step (2), the combined information—comprising of the information from the initial $00$ and $10$ states—is erased into the $01$ state; in step (3), the information in the $11$ state is translated into the $10$ state using the dynamics of the CE protocol. Note that in this step there is no logical erasure, since quadrant III was vacated in step (1). At this point, the $\mathtt{Y}^{\prime}$ output bit represents a NAND gate of $\mathtt{X}$ and $\mathtt{Y}$. Continuing to track the arrows for steps (4)-(5) completes computation $C$, assuming that $\mathtt{X}^{\prime}$ also yields a NAND gate.

### v2

@@ V.1 Effective CE Protocol
Figure 5(a) highlights the important fixed points on the potential immediately before and after the stable (red) and unstable (green) fixed points annihilate in the negative-$x$ half plane. After the annihilation, a CE is possible since the orange, blue, and purple fixed points on the positive-$x$ half plane are preserved. From the perspective detailed in Sec. III, the information stored in the $\mathtt{Y}$ state on the negative-$x$ half plane is erased, while all other information in the potential’s memory states is preserved. In this example, $\Delta U_{00}^{\prime}$ denotes the potential barrier that must vanish for the CE to be carried out. Once this barrier vanishes, $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, and $\Delta U_{11}^{\prime}$ denote the barriers that need to be maintained. The goal is to find the circuit parameter values that give these barriers sufficiently large heights, as this further contributes to a robust CE.
@@
Figure 5(b) illustrates bifurcation diagrams of the $\varphi_{2}$ coordinate over a given circuit parameter range, while all other parameters are held constant. Note that the coordinate $\varphi_{2}$ is solely illustrated because $\varphi_{1}$ varies negligibly in comparison. Additionally, if fixed points annihilate within a parameter value range, they are displayed in Fig. 5(b). Otherwise, the fixed points do not annihilate within this selected range, and their $\varphi_{2}$ trajectory will not be shown.
To guide the exploration of viable parameter values, a general view of the relationship between the dimensionless barrier heights illustrated in Fig. 5(b) and the energy scale of thermal fluctuations $k_{B}T$ is useful. Understanding this aids in determining if a barrier height is energetically large enough to store microstates within a metastable memory state.
For example, suppose the barrier should be no smaller than $50k_{B}T$. Using common values from the SQUID literature [17, 36], $T=100\;\text{mK}$ and $L=230\;\text{pH}$, this corresponds to a dimensionless barrier height of $\Delta U^{\prime}=50k_{B}T/U_{0}=0.15$. We indicate this barrier height with the dashed horizontal line in the insets of Fig. 5(b); it represents a proxy for characterizing dimensionless barrier heights. That is, if a circuit parameter’s value results in either $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, or $\Delta U_{11}^{\prime}$ falling below this line, then we say it is no longer able to reliably store information in the relevant well. Of course, this all varies depending on the device’s operating temperature and parameter values. For instance, if we wanted to continue using $U^{\prime}=0.15$ as a standard value from the QFP literature [37, 31, 38, 4]—that being $T=4.2\;\text{K}$ and $L=10\;\text{pH}$—this corresponds to a dimension-full barrier height of $\Delta U=0.15U_{0}=28k_{B}T$. As a final note, motivated by Ozfidan et al [36], we let $\beta=2.3$.
@@
The following now describes the qualitative consequences of changing each circuit parameter and the resulting potential in Fig. 5(a). First, as $\mu$ increases, the depth of the minima representing the states $01$ and $10$ increases. Meanwhile, the minima’s depth in the $11$ state decrease, leading to $\Delta U_{11}^{\prime}$ decreasing. This indicates that $\mu$ should not take on too large of a value in order to avoid fluctuations over the $\Delta U_{11}^{\prime}$ barrier. Next, if the magnitude of $\varphi_{2\mathrm{xdc}}$ increases, then both $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$ decrease. Consequently, its value should be large enough only to ensure that $\Delta U_{00}^{\prime}$ is eliminated while $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$ are maintained. Changing $\varphi_{1\mathrm{x}}$ adjusts the potential’s tilt with respect to the $\varphi_{1}$ axis. If $\varphi_{1\mathrm{x}}$ takes on too large of a value, then the barrier $\Delta U_{01}^{\prime}$ will be undesirably eliminated. However, its value must be large enough to ensure that $\Delta U_{00}^{\prime}$ is eliminated. Similarly, tuning $\varphi_{2\mathrm{x}}$ changes the tilt of the potential with respect to the $\varphi_{2}$ axis. If $\varphi_{2\mathrm{x}}$ is too large, then the blue (purple) stable (unstable) fixed points can annihilate, which causes information to be erased in the region where it should be preserved.
@@
Obtaining effective circuit parameter ranges involves determining two particular values: (i) the value resulting in $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, and $\Delta U_{11}^{\prime}$ having the same height, and (ii) the value which results in one barrier falling below the proxy line. The first (second) criterion serves as the lower (upper) end of the range. Ultimately, there are clear tradeoffs that depend on the value of a given circuit parameter. Finding the set of parameters permitting control over which information is erased and preserved requires a balancing act: A careful process of selecting parameter values that are large enough to erase the desired information, but not so large that the information originally planned to be preserved ends up being erased.
@@
Figure 5: Circuit parameter ranges that eliminate the dimensionless potential barriers $\Delta U_{00}^{\prime}$ and maintain $\Delta U_{11}^{\prime}$ and $\Delta U_{10}^{\prime}$. In the insets, light blue, light green, and light red correspond to the barriers $\Delta U_{11}^{\prime}$, $\Delta U_{10}^{\prime}$ and $\Delta U_{01}^{\prime}$, respectively. The dotted line corresponds to a thermal energy barrier of $50k_{B}T$ at $T=100$ mK with $L=230\;\text{pH}$. For exemplary purposes, $\beta$ = 2.3. (a) Specific fixed points (colored diamonds) within the potential energy landscape, as well as corresponding energy barriers, are utilized to investigate the CE in Fig. 2. (Left) Prior to the fixed point annihilation in the third quadrant. (Right) After fixed point annihilation that causes $\Delta U_{00}^{\prime}$ to vanish. Once all information stored in the $00$ state is erased into the $01$ state and the potential is subsequently brought back to its original configuration, the CE is completed. (b) Circuit parameter ranges for achieving the CE in Fig. 2, which can be accomplished by the annihilation of the green and red fixed points while maintaining the yellow and blue fixed points. Each circuit parameter range holds all other nonzero circuit parameters at the respective constant values: $\mu=0.06$, $\varphi_{2\mathrm{xdc}}=1.79$, $\varphi_{1\mathrm{x}}=0.61$, and $\varphi_{2\mathrm{x}}=0.10$. These circuit parameters lead to effective CE protocols and thereby robust NAND gates.
@@ V.2 Control Erase Parameter Selection
We now detail which circuit parameters—including their respective signs and magnitudes—are deployed for any of the eight CE protocols of interest. By leveraging the relationship between the potential’s dynamics and employing the CE notation introduced in Sec. III, all eight CE protocols in Fig. 8 can be implemented in a systematic way. The following explains this approach, summarizing the results in Table 2.
@@
Here, $[P]$ is the Iverson bracket of the statement $P$, and $[P]$ evaluates to 1 if $P$ is true or to 00 if $P$ is false. This notation is useful when discussing how the sign of a circuit parameter is determined. Let’s now review the four parameters involved in each CE protocol:
@@
For these parameters, $i,j\in\{1,2\}$ with $i\neq j$. To determine which $\varphi_{i\mathrm{xdc}}$ to use for a given CE, recall that we specify a CE with three parameters: $\rho$, $\sigma$, and $A$. If $\rho=\mathtt{X}\;(\mathtt{Y})$, then $i,j=2,1\;(1,2)$. As seen from Eq. (3), $\varphi_{i\mathrm{xdc}}$ appears only in the $\mathrm{cosine}$ function: Since this function is even, the sign of $\varphi_{i\mathrm{xdc}}$ is not relevant. Recall from Sections III and V.1 that $\varphi_{i\mathrm{xdc}}$ lowers energy barriers on both sides of the $i$th axis. However, the CE requires one barrier to be maintained. Thus, we need a control parameter to offset this barrier drop so that it applies to only one half of the $\varphi_{1}$$\varphi_{2}$ plane, while maintaining the other barrier. The parameter deploying this barrier offset is $\varphi_{j\mathrm{x}}$, and its sign is given by $\llbracket\sigma\neq\rho\rrbracket$.
Next, $\mu$ biases the potential such that the minima along one of the diagonals of the $\varphi_{1}$$\varphi_{2}$ plane  are more energetically favorable compared to those on the other diagonal. Its sign determines which well becomes the target for the erasure and is given by $\llbracket[\sigma=\rho]\ \mathrm{XOR}\ A\rrbracket$. Since $\mu$ acts along the diagonals of the $\varphi_{1}$$\varphi_{2}$ plane, it causes asymmetry between two wells that are supposed to be storing information, resulting unwanted information leakage. Fortunately, the parameter $\varphi_{i\mathrm{x}}$ can be used to compensate this effect by tilting the potential to offset this unintentional information leakage. The sign of this parameter is found through $\llbracket A\rrbracket$.
@@ V.3 CE Robustness
As this information is erased, other information is being stored in the $01$, $10$, and $11$ states due to the barriers $\Delta U_{01}^{\prime}$, $\Delta U_{10}^{\prime}$, and $\Delta U_{11}^{\prime}$ being maintained. However, we see in Fig. 5(b) that as a circuit parameter increases, at least one barrier’s height decreases. The microstates stored in these decreasing barriers are most susceptible to fluctuating out of their respective memory states. Due to this, the timescale that characterizes how long these states reliably store information is in competition with $t_{\mathrm{CE}}$. This type of timescale is known as a dwell time [16, 18, 39], which is denoted $t_{\mathrm{d}}$. Implicitly from Fig. 5(a), there are three dwell times that contribute to CE performance—those being the dwell times corresponding to the states $01$, $10$, and $11$. The minimum of these three dwell times ultimately dictates CE failure.
@@
If the length of time that information can be reliably stored is much longer than the timescale of a CE protocol, then the probability of an error resulting from an unwanted thermally activated transition is low. To quantify this, we take the ratio $t_{\mathrm{d}}/t_{\mathrm{CE}}$ as a measure of robustness.
@@
To calculate $t_{\mathrm{CE}}$, we approximate the information in the $00$ state as a single damped particle travelling down a quadratic potential from the point where the fixed points initially annihilate; the derivation is included in App. A. For $t_{\mathrm{d}}$, since there are three dwell times that affect CE performance, we choose the minimum of these dwell times to characterize CE robustness. This provides a more conservative estimate of a particular CE’s performance; refer to App. B for details. Of particular relevance, $t_{\mathrm{d}}$ is exponentially proportional to $1/L$ and $1/T$; this has a dominant effect on the robustness $t_{\mathrm{d}}/t_{\mathrm{CE}}$. Studying the robustness as a function of $L$ and $T$ allows for interpreting the performance of a CE across the MQP and QFP literature.
@@
From Sec. V.1, selecting parameter values at the lower end of the effective parameter range yields a more robust CE: We found that $\mu=0.06$, $\varphi_{i\mathrm{xdc}}=1.79$, $\varphi_{j\mathrm{x}}=0.61$, and $\varphi_{i\mathrm{x}}=0.10$. These will be known as the effective parameter values. With these, Fig. 6 displays the CE robustness for different values of $L$ and $T$. Since the MQP (QFP) literatures employ SQUIDs for different purposes—thereby leading to differing ranges of $L$ and $T$—we separate the possible value ranges into the left (right) figures. The relevant MQP literature [40, 17, 16, 18, 32, 19, 33, 27, 35, 29] corresponds to the left figure, which has a general temperature range of $[100,500]\;\text{mK}$, while the inductance values range from $[140,300]\;\text{pH}$. An upper value of $\mathcal{O}(10^{48})$ CEs per device is obtained with the inductance value of $L=140\;\text{pH}$ found from Saira et al [27] at $T=100\;\text{mK}$.
@@
Now, suppose all circuit parameters were initialized to zero, so that the potential matches Fig. 1. Then, tune all circuit parameters infinitely fast to their respective final values. Due to this, the information stored in the $00$ state is now travelling from the location determined by zero-valued circuit parameters, as opposed to the point of annihilation. This means that the information must now travel a greater distance to the $01$ state, thereby resulting in a larger $t_{\mathrm{CE}}$. Considering this scenario provides a more conservative estimate for the ratio $t_{\mathrm{d}}/t_{\mathrm{CE}}$. That said, we found that this larger $t_{\mathrm{CE}}$ will be no more than twice the value of the most conservative estimate of $t_{\mathrm{CE}}$, leading to a reduction of the robustness of order one compared to that shown in Fig. 6.
@@ V.4 NAND Gate Application
Executing successive CE protocols allows for performing a NAND gate. The CE has two input bits and two output bits as seen from Fig. 3, while the NAND gate has two inputs but only one output bit. When the number of logical output bits is less than the number of embedded output bits in a computational system, there is added flexibility when choosing which bit to read as the output for a computation. For example, we can specify $\mathtt{X}^{\prime}=\text{NAND}[(\mathtt{X},\mathtt{Y})]$ without prescribing the $\mathtt{Y}^{\prime}$ computation. Thus, we define two kinds of NAND gates: a partial NAND, written as $P_{\rho^{\prime}}[(\mathtt{X},\mathtt{Y})]$, which only requires the state $\rho^{\prime}$ to yield the desired output, and a complete NAND, denoted as $C[(\mathtt{X},\mathtt{Y})]$, which requires that $\mathtt{X}^{\prime}=\mathtt{Y}^{\prime}=\text{NAND}[(\mathtt{X},\mathtt{Y})]$. Fig. 7(a) tabulates these choices.
From here, Fig. 7(b) illustrates both kinds of computations, and Fig. 7(c) tabulates the computations. Recall from Fig. 1 that the first, second, third, and fourth quadrants correspond to the memory states $11$, $01$, $00$, and $10$, respectively. In this example, carrying out only steps (1)-(3) executes the partial computation $P_{\mathtt{Y}^{\prime}}$. This can seen by tracking the arrows through the included protocol steps: In step (1), the information in the state $10$ is erased into the $00$ state; in step (2), the combined information—comprising of the information from the initial $00$ and $10$ states—is erased into the $01$ state; in step (3), the information in the $11$ state is translated into the $10$ state using the dynamics of the CE protocol. Note that in this step there is no logical erasure since quadrant III was already vacated in step (1). At this point, the $\mathtt{Y}^{\prime}$ output bit represents a NAND gate of $\mathtt{X}$ and $\mathtt{Y}$. Continuing to track the arrows for steps (4)-(5) executes the computation $C$.